Of the $50,000 that Natasha pocketed on her last real estate deal, $20,000 went to charity. She invested part of the remainder in leaders Fund with an annual yield of 16% and the rest in another Fund with an annual yield of 25%. If she made $6060 on these investments in 1 year, then how much did she invest in each fund?
Answer by Earlsdon(6098)

Natasha invested $30,000 ($50,000-$20,000). Let x = the amount she invested at 16%, then ($30,000-x) would be the amount she invested at 25%. The interest earned on these two amounts can be expressed as:
0.16x+0.25($30,000-x) = $6,060 Simplify and solve for x.
0.16x+7500-0.25x = 6060 Combine the x-terms.
-0.09x+7500 = 6060 Subtract 7500 from both sides.
-0.09x = -1440 Divide both sides by -0.09
x = $16,000 and...
$30,000-x = $14,000
Natasha invested $16,000 in Leaders Fund at 16% interest and $14,000 in another Fund at 25% annual interest.